Five years ago, Tom was one third as old as his father was then. In 5 years, Tom will be half as old as his father will be then. Find their ages now.
Answer by scott8148(6628)

"Five years ago, Tom was one third as old as his father was then" ___ 3(t - 5) = f - 5 ___ 3t - 15 = f - 5 ___ 3t - 10 = f

"In 5 years, Tom will be half as old as his father will be then" ___ 2(t + 5) = f + 5 ___ 2t + 10 = f + 5 ___ 2t + 5 = f

substituting ___ 2t + 5 = 3t - 10 ___ 15 = t

substituting ___ 2(15) + 5 = f ___ 35 = f